### v2

# Deciphering the global production network from cross-border firm transactions

Abstract
Critical for policy-making and business operations, the study of global supply chains has been severely hampered by a lack of detailed data. Here we harness global firm-level transaction data covering 20m global firms, and 1 billion cross-border transactions, to infer key inputs for over 1200 products. Transforming this data to a directed network, we find that products are clustered into three large groups including textiles, chemicals and food, and machinery and metals. European industrial nations and China dominate critical intermediate products in the network such as metals, common components and tools, while industrial complexity is correlated with embeddedness in densely connected supply chains. To validate the network, we find structural similarities with two alternative product networks, one generated via LLM queries and the other derived by NAFTA to track product origins. We further detect linkages between products identified in manually mapped single sector supply chains, including electric vehicle batteries and semi-conductors. Finally, metrics derived from network structure capturing both forward and backward linkages are able to predict country-product diversification patterns with high accuracy.
Keywords: Supply chains, resilience, policy, network science
Disclaimer: This paper contains the views of the authors and not necessarily those of the UK Government or its Ministers.
Acknowledgements: We would like to thank Christian Diem and Tobias Reisch for their useful input and comments, and Shaun Hoang for his research assistance. The data and computing infrastructure was provided by the the UK Government Global Supply Chain Intelligence Programme (GSCIP). NOC would like to acknowledge funding from the 2024-2025 UKRI Policy Fellowship programme.
Introduction
As acutely highlighted by the global pandemic, limited insight into global supply chains can have serious and immediate consequences. However, until quite recently, supply chain analysis was limited to specific sectors or highly aggregate economy-wide transaction data which lacks detail and hence is not very useful or sufficiently accurate for many policy and research questions. Here we harness ’next generation’ global inter-firm transaction data to infer detailed product inputs as the basis for constructing a product level supply chain network.
Traditionally, flows and transactions between sectors in the economy have been captured via input-output tables (Leontief, 1986). These capture monetary flows between firms within defined industries, and are typically produced by national statistical agencies. Some countries, such as the US and Mexico, produce these at quite a detailed level (4 digit NAICS - or approx 400 industries). However, in most countries these are only available for very aggregate sectors (in the UK, for example, IO tables cover just 105 sectors). And, importantly, they don’t capture cross border flows limiting their potential for global supply chain analysis. There have been a handful of efforts to build integrated input-output networks that include inter-country flows, but again these are extremely aggregate in nature. For example, the World Input-Output Dataset (WIOD) (Timmer et al., 2012) has just 56 sectors and 40 countries (2010-14) or the OECD World Input-Output Table (WIOT) with 36 sectors and 69 countries (1995-2020). A variety of studies have investigated the structure of the World Input-Output Network (WION), finding strong geographical clustering of sectors within neighbouring countries, but also some instances of strong inter-country clustering in sectors such as textiles and transport equipment Piccardi et al. (2018).
In recent years there has been a huge push to better collect and make available for research firm level transaction data (Pichler et al., 2023). Research using this data has indicated that, for example, the propagation of shocks at the firm level has distinct characteristics and magnitudes which cannot be approximated by aggregate data (Diem et al., 2022, 2024; Inoue and Todo, 2024). Up until very recently, however, most firm level data was collected by single countries (e.g. VAT data) or for a specific purpose (i.e., shipping data), and hence had very limited coverage.
At the other end of the spectrum are efforts to manually reconstruct detailed supply chains for specific sectors. These studies collect data from a variety of sources including internal systems like procurement, inventory management, and logistics, as well as external sources such as supplier data and transportation information, to build a detailed picture of the key inputs needed to produce often strategically important goods. For example, there have been careful efforts to map out antibiotic (Klimek et al., 2023), semi-conductor (Haramboure et al., 2023) and battery supply chains (Caravella et al., 2024). However, these supply chain mapping exercises tend to focus on material inputs - critical minerals and chemicals - rather than machinery, tools and other necessary inputs such as cleaning equipment. And, critically, their fragmented nature makes them not suited to many questions which require a wider structural lens on product inter-connectivity.
A major focus of analytical efforts with respect to supply chains has been to been to better understand complex global inter-dependence structures and vulnerability to shocks. Risks might come from geographic shocks, such as localised shipping disruptions or natural disasters, sector specific shocks, perhaps due to tariffs or missing critical inputs, or broad based shocks such as the pandemic or financial crisis. Much of this literature aims to quantify node importance - whether it be nodes that influence the degree of spreading processes, or those that are most susceptible to impacts. A number of papers suggest that the US IO network follows a power law structure (Acemoglu et al., 2012; Han and Goetz, 2019; Chopra and Khanna, 2015), and that a small number of key sectors (including critical infrastructure sectors) are involved in many economic transactions and are hence susceptible to inducing large scale cascades. Blöchl et al. (2011) proposed two centrality measures to capture the impact of shocks: random walk centrality to capture most immediate impacts, and betweenness to capture where a shock lingers longest. The authors apply these metrics to the OECD WIOT, finding ’wholesale and retail trade’ is most often the node with highest random walk centrality while ’motor vehicles’ have the highest betweenness.
Supply chain linkages are also key to longer term impacts such as diversification of countries and regions into new economic activities, otherwise referred to as ’upgrading’ (Kaplinsky, 2000). Originally proposed as a mechanism to aid countries with natural resources diversify into productive activities (Hirschman, 1958; Gereffi et al., 1999), it is widely understood today that embeddedness in global supply chains aids both domestic productivity and export diversification (Pahl and Timmer, 2020). While a wide array of mechanisms have been proposed, many focus on accumulation of technological capabilities and the diffusion of knowhow, as espoused by the Evolutionary Economic Geography (Nelson and Winter, 2002; Boschma and Frenken, 2006) and Innovation Systems (Edquist, 2013) literatures. From this perspective, supply chain linkages are thought to drive knowledge diffusion via, for example, demands for quality upgrading, worker flows and increased competition (Blomström and Kokko, 1998; Criscuolo et al., 2017). These effects are thought to be particularly strong in the case of supplier-customer relationships between local firms and co-located foreign multi-nationals (Javorcik, 2004), but there are greater uncertainties about the potential for such linkages to induce upgrading when it comes to linkages spanning large geographical and technological distances (Pietrobelli and Rabellotti, 2011).
The empirical evidence on the benefits of participation in global value chains is mixed (as reviewed by Kowalski et al. (2015)), finding a strong influence on manufacturing productivity and export sophistication, but mixed impact, for example, on domestic employment (Pahl and Timmer, 2020; Gonzalez, 2016). In terms of direction, both backward linkages (i.e., diversification into more complex downstream goods) and forward linkages (i.e., into input or supplier goods) have been proposed as channels for diversification. Methodologically closest to our work, Bahar et al. (2019) find evidence for the role of forward linkages in international export product diversification patterns over a 10 year period. Linkages in this case, however, were proxied from US IO data measured for US domestic industrial sectors and do not capture global value chain relations.
Beyond the short-medium term impacts of shocks, a recent literature studies the influence of production network structure on local economic growth and resilience. Taking an evolutionary stance, in this context resilience is typically defined as the extent to which regions or countries adapt to shocks in the longer term rather than return to pre-shock norms (Boschma, 2015). Empirical evidence at a subnational level supports the idea that regions’ diversification patterns are conditioned on their ’role’ in national production networks. For example, US cities tend to enter new industries that are ’related’ to those that already exist (Essletzbichler, 2017), where relatedness is measured in terms of supply chain proximity (Fan and Lang, 2000). When it comes to resilience, Han and Goetz (2019) show that weighted centrality metrics derived from the US IO network are associated with employment resilience for US counties during the 2007-2009 recession, while Cainelli et al. (2019) find that EU regions with highly integrated production networks were worse off during the same period (which the authors suggest is due to enhanced amplification of the shock).
Here we aim to infer detailed ’product recipes’, or the set of unique inputs needed in the production of a specific product. In other words, we aim to capture the input materials and tools used in product that are specific to a particular product, and not common across many products (e.g., packaging). Distinct from two recent efforts build a similar product network using firm-level transaction data from Italy (Fessina et al., 2025) and Japan (Inoue and Todo, 2024), we aim to capture global supply chains (as opposed to local supply chains) via global firm purchasing patterns. To do this, we harness ’next generation’ global firm-level transaction data via a UK government programme, GSCIP. We use an aggregated subset of data from GSCIP for the years 2021-2023, prior to additional augmentation of the transactional data completed in 2025. The subset of data that we use is orders of magnitude larger than single country datasets typically in use today, and covers 20m global firms, and 1 billion cross-border transactions. The data was assembled from a variety of sources, from customs to firm-owned data, and matched at a firm level.
To detect the set of inputs required to make a specific product, we compute the excess purchase of inputs (products) for the set of firms that make a specific product relative to how often an average firm purchases the input. This latter step effectively down-weights common or ubiquitous inputs. This approach echoes that of O’Clery et al. (2021), in which the authors inferred country-product diversification paths by extracting the set of common products produced in the set of countries in which a specific new product appeared later. In this case, one can think of this approach as ’piecemeal’ reconstruction of global product supply chains in that we independently estimate each set of input linkages for each product. In this way, we capture both inputs or ingredients, and complex tools and packaging.
These relationships can be represented and analysed as a network in which nodes are products, and directed weighted edges represent the relative importance of the input. Analysis of the meso-structure of this network suggests that textiles’ supply chains are largely isolated from wider manufacturing, while chemicals are typically more inter-connected with food supply chains than machinery and metals. European industrial nations and China dominate nodes with high betweenness centrality, i.e., critical intermediate products in the network, while countries with high industrial complexity (proxied via the ECI metric of Hidalgo and Hausmann (2009)) tend to export goods that have a high ’hub score’ in the network, i.e., they are embedded in densely connected supply chains.
While there is no ‘ground truth’ data available, we have undertaken significant efforts to validate the network structure. We found structural similarities with an alternative AI-generated product network (Fetzer et al., 2024), and a product supply chain dataset created by NAFTA to track product origins (Conconi et al., 2018). In addition, we found dense connectivity in our network between nodes identified as belonging to two manually constructed supply chains - electric vehicle batteries and semi-conductors. Finally, metrics derived from the network structure are able to predict country-product diversification patterns with high accuracy, finding evidence for both forward and backward linkages.
Results
We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. Firm level transaction data has been compiled from a variety of sources including customs and firm-held data, such as shipping data. These sources have been integrated using AI-based tools to, for example, disambiguate firm names and match them. The data we used represents a static cut which was made available to the research team via a UK government programme (GSCIP), includes firm level attributes for over 20m firms and 1 billion inter-firm transactions between 2021-23. All transactions are between firms located in different countries. The GSCIP data contains additional firm-level information from other sources, such as firm ownership structures.
We have a number of firm and transaction attributes, a selection of which is shown in Figure 1 B, including dollar transactions for specific products between firm pairs. As shown in Figure 1 C-D, the data has high coverage across developed economies, and those with high trade (exports plus imports) volumes such as China and Vietnam. It also has good coverage in some middle income countries such as India and Peru, but with lower coverage overall in less economically active regions such as Africa.
Figure 1: We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. [A] Firm level transaction data has been compiled from a variety of sources. [B] This static cut of GSCIP data includes firm level attributes for 20m firms, combined with 1 billion inter-firm transactions. [C-D] The data has very good coverage of firms in developed economies, with lower coverage in less economically active regions such as Africa. Top 50 countries by firm count shown.
We infer the importance of each product as an input to any other product from the firm transaction data. We aim to derive a metric to capture the ’importance’ of product $i$ for the production of product $j$. In essence, we compute the ’excess’ purchase of product $i$ by firms that produce product $j$, relative to the purchase of $i$ by ’an average firm’. In our toy example in Figure 2 B, we show that while aluminium panels are purchased very frequently by washing machine firms (relative to all firms), cardboard boxes are often purchased by all types of firms including washing machine firms and hence they are a very common input (and are thus down-weighted in our scheme).
More precisely, we set stringent definitions to identify firms that both produce and purchase specific products:
  • A firm $k$ produces product $j$ if they sell a larger monetary value of $j$ than the average firm that sells product $j$.
  • A firm $k$ buys product $i$ if they purchase a larger monetary value of $i$ than the average firm that buys product $i$.
The set of all firms is denoted $S$. For each product $j$, we form a subset of firms that produce product $j$, denoted $S_{j}$. A subset of these buy product $i$, $S^{i}_{j}$. Finally, $S^{i}$ is the set of firms that purchase product $i$ (irrespective of what they produce).
Our ratio is the share of firms who produce $j$ that also buy $i$ divided by the overall share of firms that buy $i$:
$A_{i,j}=\frac{|S^{i}_{j}|/|S_{j}|}{|S^{i}|/|S|}$
If $A_{i,j}>1$, then product $i$ is more prevalent in the purchasing basket of firms that produce $j$ than the wider set of all firms. We can also set a minimum for the size of the set $|S^{i}_{j}|$ (the number of firms that both produce $j$ and buy $i$), which we refer to as parameter ’firmcount’. In essence, for low values of firmcount, we have few ’pieces of evidence’ for the edge weight (and so for some applications we might want to set a higher firmcount threshold, which would make our network sparser). In the analysis below, we often show results for a range of values of both the edge weight threshold and firmcount.
As shown in Figure 2 B, we pick up both inputs or ingredients and complex tools. For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
Figure 2: We infer the importance of each product as an input to any other product from the firm transaction data. [A] We compute the excess purchase of product $i$ by firms that produce product $j$, relative to the purchase of $i$ by ’an average firm’. In our toy example here we show that while aluminium panels are purchased very frequently by washing machine firms relative to all firms, cardboard boxes are purchases by all types of firms including washing machine firms and hence they are a very ubiquitous input (and are thus down-weighted in our scheme). [B] We pick up both inputs or ingredients and complex tools. For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
We transform the product input relations into a network, and find three main clusters representing machinery and metals, chemicals and food, and textiles. We transform the product input relations into a directed weighted network, with edge weight stored in adjacency matrix $A$ (with entries $A_{i,j}$ as above) of size $n$ x $n$ for $n=1228$. In Figure 3 A, we observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firmcount parameter, the edges become more sparse, but the edge weight threshold has no major effect. In practice, the majority of the analysis to follow focuses on the largest weakly connected component of the network with edge threshold value 2 and firmcount value 2, with size $n=1207$. In many cases, and also in the SI, we also show results when varying these parameters.
This network can be visualised using a ’spring’ based visualisation algorithm (Force Atlas in Gephi was used here), as shown in Figure 3 B. We observe loose clustering of nodes according to HS ’Section’, a high level classification of products into sectors. However, rather than rely on an external classification which bears no relation to the inter-connectivity we observe in our data, we use a multi-scale community detection algorithm to uncover the ’meso’ structure or high-level organisation of the network into node ’communities’. These are clusters of products with strong internal connectivity - but relatively weaker connectivity to the rest of the network. These structural patterns are highly informative, capturing for example the behaviour of cascades and shock propagation (Fortunato, 2010).
While there are many distinct approaches to community detection suited to different applications (Fortunato, 2010), in our case, we employ a multi-scale algorithm based on random walkers on a network (Delvenne et al., 2010). Essentially, as the walkers roam on the network, jumping from node to node based on probabilities derived from the directed edge weights, we detect larger and larger clusters with strong internal connectivity. Hence, as ’time’ goes on, the algorithm detects fewer but larger communities. We can learn from this process: as the algorithm effectively ’forces’ communities to merge into larger clusters, the ’time’ at which they merge (or refuse to merge) is highly informative about their relative inter-connectivity. In other words, those that merge earlier are more connected, but those that remain unmerged for longer times are more isolated and disconnected from the wider network.
Here, shown in Figure 3 C-E, we show three network partitions into communities at three distinct ’times’ (see SI for more details and further analysis). The merging process suggests that textiles is highly isolated in the sense that it remains unmerged across all three partitions shown. We find that initially (Figure 3 C) machinery is split across two clusters - one focused on electrical goods and the other on metal-based tools (and metals themselves). We observe that over time, food merges with plastics and then chemicals, and both machinery clusters merge. Then, these two groupings merge - but textiles and some other small clusters (musical products, watches, and art) remain unmerged signifying their isolation. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
Figure 3: We transform the product input relations into a network. [A] The product supply chain network can represented by an adjacency matrix in which a value in entry $A_{i,j}$ denotes a non-zero edge weight between node (product) $i$ and node $j$. We observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firm-count parameter, the edges become more sparse, but the edge weight threshold has no major effect. [B] The network can also be visualised using a visualisation algorithm (Force Atlas was used here). [C-D] We use a multi-scale community detection algorithm to uncover the ’meso’ structure of the network. As the algorithm effectively ’forces’ communities to merge into larger clusters, we find that textiles is highly isolated in the sense that it remains unmerged over many scales. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. In the SI we show that machinery products including those related to cars and motors tend to have a large in-degree (they require many inputs), while components, plastics and packaging tend to be inputs to many products. Further, consistent with what we might expect, deploying a UN classification that sorts products into capital, intermediate and consumption goods, we find that capital products are inputs to many products (high out-degree), while both capital and consumption goods have high in-degree – they require many ingredients.
Betweenness centrality (BC) measures the extent to which a node acts as a key ’juncture’ through which many paths (i.e., supply chains) cross. Mathematically, the betweenness centrality of a node is the number of times a path (between any pair of nodes) paths transits that node. In Figure 4 A, we highlight nodes with high betweenness centrality, including various types of machinery, common components such as pumps, blades, taps and motors, plastics, packaging, and chemical binders. Nations which produce these key inputs have an important role in global supply chains, controlling so-called ’choke points’. To investigate this, we compute the Revealed Comparative Advantage (a measure of export concentration of a country in a product) of 113 countries from global trade data (Comtrade 2024). In Figure 4 B, we show a global map with countries shaded by their share of top-50 BC products exported with RCA. We find European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
Highly industrialised nations engage in complex production processes via embeddedness global supply chains. To capture production complexity, we adopt an approach proposed by Hidalgo and Hausmann (2009) to infer product complexity (PCI) from patterns of geographical production. To measure the complexity of the supply chain in the vicinity of a product in our network, we compute the hub score of each node. This is similar to eigencentrality in the sense that it captures ’neighbour of neighbour’ linkages, but is suitable for directed networks such as ours. We find that the PCI (obtained for 2023 from the Harvard Atlas of Economic Complexity) has a rank correlation of 0.46 with node hub score. In essence, complex products are embedded in complex supply chains. At the country level, the country-level Economic Complexity Index (ECI), which is mathematically equivalent to the mean PCI of products exported with RCA, has a high rank correlation of 0.85 with mean hub score. Hence, countries with high industrial complexity tend to export products which are highly embedded in densely connected global supply chains.
Figure 4: European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. [A] Various types of machinery, common components such as pumps, blades, taps and other common ’ingredients’ have high betweenness centrality. [B] European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products. [C] Node ‘hub score’ is similar to eigencentrality in in the sense that it captures ’neighbour of neighbour’ linkages. [D] The Product Complexity Index (rank) correlates 0.46 with hub score. [E] Economic Complexity Index (ECI) has a high rank correlation of 0.72 with mean hub score. Hence, countries with high industrial complexity tend to export products which are embedded in densely connected global supply chains.
The supply chain network correlates with other efforts to reconstruct inter-product linkages. First, we consider a recent effect to construct a similar type of product network using LLM queries (Fetzer et al., 2024). In essence, the authors asked an LLM what are the top input products for each product, and constructed a network based on its answers in a similar fashion to above.
Figure 5 A shows that the (binary unweighted) AI network is much sparser than our network (see adjacency matrix in inset). Its clear from the distribution of white nodes (those with zero in-degree) that the AI network tends to have much fewer edges in sectors such as chemicals, plastics and food. Overall, our network (the SC network) has 4 times as many links as the AI network, but 5.8 times the number of links in the SC food, 4.5 times the number of links in the SC chemicals community and 3.3 times the number of links in the metals and electric machinery community. We compare communities for the AI and SC network in Figure 5 C. We find that the AI and SC product networks have similar meso-scale structure, with the exception of much larger and stronger chemicals and food clusters in the SC network.
Next, we compute a simple correlation of the full edge set of AI network and SC network, finding they are correlated up to approx 0.35 for low values of firmcount (i.e., including uncertain edges). The in- and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree. In general, these correlations peak for low firmcount in the range 2-4 - i.e., we keep weak edges for a dense network. Hence, while the networks look visually very different, they display non-trivial similarity in both meso structure and in- and out-degree pattern.
We also find similarity to a NAFTA derived product-level IO network. We compare our SC network to a NAFTA derived product-level ’IO network’ from Conconi et al. (2018) which was constructed based on NAFTA ’Rules of Origin’ which aims to identify the origin country of goods in order to determine if tariffs apply. In this case we find a much lower edge correlation of 0.15, a binary in-degree correlation of 0.28 and out-degree correlation of 0.32. Despite the lower correlations, however, the aggregate meso structure retains similarities to the SC (and AI) networks with food, textiles, chemicals and metals/machinery clusters (see Figure S3).
Figure 5: The product network correlates with other efforts to construct inter-product links or chains. [A] The ’AI network’ of Fetzer et al. (2024) is constructed via LLM queries. Nodes with zero in-degree are shown in white. [B] The AI network is much sparser than the SC network and tends to have much fewer edges in sectors such as chemicals, plastics and food. [C] The AI and product network have similar meso-scale structure, with the exception of a much larger and stronger chemicals and food clusters in the SC network. [D-E] The correlation of the full edge set of AI network and SC network is approx 0.34, while the in and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree (for low firmcount in the range 2-4).
The product network picks up links identified in manually constructed single-sector supply chains. Up until now, most detailed data on inter-product linkages has come from a large number of detailed manual studies of sector specific supply chains. Here we take a selection of these, and investigate to what extent we pick up similar linkages using our method. We focus on two key sectors, electric vehicle batteries and semiconductors. For each sector, we take the set of products identified in the manual supply chains, and visualise the entries of the adjacency matrix of the subgraph of these products in Figure 1 A/C. In other words, we show the edge pattern that we find using our method for the identified product sets. We observe that we capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. While products in electric vehicle battery are highly clustered in the centre of the network, semi-conductors exhibit two distinct clusters composed of chemicals and machinery.
We wish to show that statistically its very unlikely we could have found these links at random. To do this, we simulate 100k random networks with the same number of nodes (as those identified in the sector supply chain) and compare the modularity score of our subgraph to those of the random networks. The modularity score essentially tells us how modular or connected the subgraph is compared to what would be expected at random under a null model (we use a standard null model which randomly allocates edges while keeping the node degree fixed). We compare the modularity of our network (far right axis) to the distribution of modularity scores for the ensemble of 100k subgraphs, finding an extremely low probability that our subgraph would be simulated at random (pvalue shown).
Figure 6: The product network picks up links identified in manually constructed single-sector supply chains. [A/D] We capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. [B/E] While products in electric vehicle battery are highly clustered in the centre, semi-conductors exhibit two distinct clusters composed of chemicals and machinery. [C/F] To show we pick up linkages between these products that are far from random, we simulate 100k random networks with the same number of nodes and compare the modularity score (connectivity) of our subgraph (see A/D) to the random networks. We observe that our subgraphs are in the very extremes of the distributions suggesting they would be extremely unlikely to occur at random.
Both downstream and upstream links predict country-product export diversification patterns. Finally, we probe whether metrics derived from our network are predictive of country-product diversification patterns. To do this, we use a similar approach to a large literature on density metrics in EEG (Hidalgo et al., 2018) to construct a ’downstream’ (and analogous ’upstream’) density metric. As illustrated in Figure 7 A/D, essentially, for a product $p$, this captures the share of top in-degree neighbouring products present in country $c$:
$d_{p,c}=\frac{\sum_{j\in J_{p}}I(A_{p,j})M_{p,c}}{\sum_{j\in J_{p}}I(A_{p,j})}$
where $J_{p}$ is the set of top $k=50$ edges from $p$ to all nodes, and $M_{p,c}=1$ if product $p$ is present in country $c$ (and otherwise 0). We construct an equivalent ’upstream density’ metric.
Similar to O’Clery et al. (2021), we perform a standard Probit regression for the probability of a product appearance of the form:
$R_{p,c}=\Phi(\alpha+\beta_{d}d_{p,c}+\gamma_{p}+\eta_{c})$
where $R_{p,c}=1$ if product $p$ was present in country $c$ in 2021 but absent in 2016, $\Phi$ is a normal cumulative distribution function, and $\gamma_{p}$ and $\eta_{c}$ are product and country fixed effects respectively.
The tables in Figure B/E show that both downstream and upstream links predict country-product export diversification patterns with high predictive power (the AUC or ’area under the curve’ measures the model’s ability to distinguish between positive and negative cases) for both models with and without fixed effects. The AUC values we find are in line with the literature on predicting product entries in countries, and other similar efforts to predict diversification on a range of scales using density metrics derived from a variety of networks capturing ’capability overlap’ between products (Hidalgo et al., 2007; O’Clery et al., 2021).
Figure C/F shows, perhaps counter-intuitively but in line with some of the validation steps above, that the predictive power is highest for low firmcount. Hence, including ’weak links’ for which there is relatively lower certainty is informative for the model. The predictive power is also higher for lower entry threshold - i.e., we better predict small increases in Rpop (which compares a country’s per capita production levels in a product to the world’s overall per capita production of the product) over a 5 year period compared to larger jumps.
Figure 7: Both downstream and upstream links predict country-product export diversification patterns. [A/D] Downstream/upstream-density captures a country’s presence in products that are downstream/upstream of the target product. [B/E] Both downstream and upstream links predict country-product export diversification patterns with fairly high predictive power (AUC) for cases with and without fixed effects. [C/F] The predictive power is highest for low firmcount (i.e., including ’weak links’), and for low rpop entry threshold (i.e., we predict small increases in rpop (product concentration relative to GDPPC).
Discussion
In this paper we develop a highly granular network tool to describe complex global inter-related product supply chains. We explore its broader structural characteristics, and investigate which countries dominate common input nodes with high betweenness, and which are active in highly embedded ’hub’ nodes. Despite an absence of ’ground truth’ data, we validate it in several ways including comparing it to a variety of other efforts to construct product level supply chain networks, and manually constructed data on sector specific supply chains.
There are a wide range of potential research and policy uses for this type of tool. From a policy viewpoint, this tool could be used to, for example, quickly map supply chains for new products which have not already been mapped manually, as well as capture inputs not typically considered in manual mapping such as machinery and tool inputs. More broadly, when augmented with country level information, the tool can be used to gain insights into the processes underlying globalisation and the ’unbundling’ of supply chains (Baldwin, 2006).
The network structure provides key insights on ’central nodes’ and clusters of highly interconnected products. Central nodes might be worthy of particular focus or protection. For example, the fact that steel has very high betweenness centrality indicates that it likely plays a role in many paths (chains) and might be seen as a critical good. This accords with the fact that the UK government moved aggressively to protect domestic production of domestic virgin steel in early 2025 in the face of impending closure of British Steel. More broadly, the network can be used to quantify supply chain vulnerabilities and model the propagation of shocks and policies such as tariffs in a more detailed, comprehensive manner than previously possible. In the case of product clusters, we can see network communities as a sort of alternative product hierarchy relative to official sectoral classifications. Highly connected clusters would naturally be collectively impacted by many kinds of cascading shocks and so they might be collectively targetted by policy interventions rather than individually.
In terms of wider applications, there is a growing literature on ’embodied carbon’ as carbon from input products is incorporated to final goods (Sato, 2014). Little is known, however, about so-called Scope 3 emissions - or those transmitted via supply chains - despite the fact they thought to be growing faster than other types of emissions (Hertwich and Wood, 2018). This type of detailed network which captures - at least in theory - the full set of product relations could be used to track ’product life-cycle’ emissions as products are manufactured and then used in other manufacturing processes (Wang and Gupta, 2011).
Here we show that country-product export diversification is associated with both forward and backward linkages, with a slightly stronger signal for backward linkages. The network could be used as a policy tool to identify opportunities in the form of upstream and downstream products which could be targets for industrial policy support. This type of policy approach is used, for example, in ’Smart Specialisation’ efforts at a European level in which regions aim to specialise and diversify in related sectors (Balland et al., 2019). Beyond nodes, it also sheds light on potential new edges. For example, it could be used to identify linkages which exist globally but not locally.
There are some limitations to our approach, as well as potential sources of noise. In order to build our network, we aggregate data over 3 years based on advice from the supplier and our analysis of the raw firm-interaction data that suggests the data is most complete for these latter years. This approach, however, does not currently allow for temporal analysis. Future work may investigate the extent to which disaggregating into distinct time periods and recomputing the network for each is feasible. Secondly, since our transactions are between entities located in different countries, we do not capture local linkages. For example, if the water used to produce beer is predominantly purchased locally, we may not capture it as a key input. Hence, we over-emphasise inputs and relationships in global value chains rather than local ones.
Third, there is no doubt some noise in our data due to the uneven composition of the firm-level data. For example, while we have good coverage of firms across developed countries, we by no means have the ’universe’ of firms in our dataset, and some countries/sectors will be more represented than others. This issue is less problematic for our tool than say country-specific or firm-level analysis, but nonetheless it will introduce noise which we cannot fully mitigate against. Fourth, there may be other sources of noise, such as short term changes in purchasing patterns. For example, we noticed an unusual input to semiconductors - washing machines. But it turned out that due to a huge shortage in input components in 2022/23, semiconductor firms were actually buying washing machines to strip them of their chips. Another source of noise is likely to be multi-product firms. While we have taken efforts to reduce this, by omitting firms that sell products in more than five HS Sections, and looking for common purchases by all firms that make a product, there will no doubt be some noise remaining due to this issue.
Fifth, there may be idiosyncratic differences between supply chain relationships within regions that we aggregate over by computing a network using all firm transactions. For example, battery supply chains may look different in the US vs Asia. While we have looked at some geographic variations here (see SI), in future work we may reconstruct the network using data for specific regions to further investigate this issue. Finally, we present the network at HS 4-digit level. It may be possible in future work to construct the network at 6-digit level, but due to an exponential increase in the number of edges to estimate, we have much fewer number of firms (firmcount parameter) for each linkage and a more complicated validation exercise with fewer sources of external data to validate against. Again, we leave this to future work.
Methods
Processing of underlying firm data
The cut of data we used from GSCIP on firm-level transactions covers the time period between Jan 2021 and Dec 2023.
We take the following processing steps:
  • We assign all firms to an ’owner firm’ based on ownership information available to us (see SI).
  • We aggregate all firms to an ’owner-country’ level. In other words, we have Coca Cola-Japan and Coca Cola-Belgium.
Construction of supply chain network
A concern about our approach might be that it could pick up spurious links due to the presence of firms in our dataset which make many products (and hence we cannot infer specific inputs to products from their purchases). Our methodology is designed to reduce this issue since we look at the purchases of not just one firm but all firms that make a specific product. To mitigate this potential issue further, we drop all firms that export products in five or more HS Sections (e.g., this would exclude wholesalers and multi-nationals with large portfolios of products such as Proctor and Gamble).
In order to further reduce noise, in addition to the firmcount parameter as mentioned above, we also drop any links from product $i$ to product $j$ in which the total value of the transactions of the firms in set $S^{i}_{j}$ is less than \$1000.
Network analysis
Community detection We use the Stability algorithm of Delvenne et al. (2010) to detect communities in the SC and AI networks. The algorithm detects partitions on a range of scales corresponding to random walker time on the network. The higher the resolution, the longer the time, the larger (and fewer) the communities. We use the ’full’ (non-linearised) specification for a directed Laplacian, with 100 iterations. In the SI we show the number of communities, variation of information across the iterations (a low value tells us we find similar partitions for different random seeds), and the stability score for our set of partitions corresponding to a range of timescales.
Modularity score We use the Modularity score to compare the connectivity of the subgraph induced by each of our sector-specific supply chains (EV batteries and semi-conductors) to 100k randomly generated subgraphs of the same size. Modularity (Newman, 2006) compares the connectivity of a community to what would be expected under a random null model - in this case a network in which the node degree is preserved but the edges randomised.
Effectively, we decompose the directed Modularity score proposed by Leicht and Newman (2008) to evaluate a set of communities for a full network, and use it instead to evaluate the connectivity of a subgraph (i.e., a single community). This is mathematically coherent as Modularity is a sum over the connectivity within communities. Specifically, we compute:
$M_{G}=\frac{1}{m}\Large(\sum_{i,j\in G}X_{i,j}-\frac{Out_{i}In_{j}}{m}\Large)$
for node pairs $i,j$ in subgraph $G$, $X=A>0$ (unweighted adjacency matrix), $m$ is the number of edges in the full network, $Out_{i}$ is the out-degree of node $i$ and $In_{j}$ is the in-degree of node $j$.
Econometric analysis
To capture product entries, we wish to compute a measure of a countries’ concentration in a product. The traditional approach is Revealed Comparative Advantage (RCA) proposed by Balassa (1965) captures is the proportion of the country’s exports in a product divided by the global proportion of exports in a product. We use a related approach, denoted Rpop, which helps remove noise due to fluctuations in commodity prices, seasonal employment or currency exchange rates (Hidalgo, 2021). Specifically, it compares a country’s per capita production levels in a product to the world’s overall per capita production of the product:
$Rpop_{p,c}=\frac{E_{p,c}/pop_{c}}{E_{p}/pop}$
where $E_{p,c}$ is the total exports of country $c$ in product $p$, $E_{p}$ is global exports of product $p$, $pop_{c}$ is the total population of country $c$ and poppop is the global total population.
We conduct the econometric analysis for a set of 235 countries. We drop products not in the main connected component of the network and those with relatively small global trade (less than 2 billion\$ in 2016) for a total of 683 products. Of a possible 235x683 observations, 86% were absent ($Rpop<0.05$) in 2016. Of these, 3.5% of products appeared by 2021 (with $Rpop>0.1$). While we keep the absence threshold fixed, we show results for a range of presence thresholds in Figure 7 C and F.
Supplementary Information
Overview of underlying data
Data sources and coverage
This paper relies on data provided by the UK HM Government’s Global Supply Chain Intelligence Programme (GSCIP). This contains firm-to-firm cross-border trade, derived from customs and shipping data, with details on the transacting firms compiled from tax filings, company websites, and other sources. Firm-to-firm transaction data is derived from administrative customs data. GSCIP amalgamates first-party customs data across a multitude of countries.
Our dataset includes transaction information for every country. While the dataset does not have first-hand customs data or another primary source of transaction information for every country, the data that we do have includes transactions between firms within the countries included and those out of sample providing partial coverage.
We are using a static cut of the data, which has further been restricted to the period from January 2021 to December 2023 as this period had the greatest coverage for the subset.
Data processing of underlying firm data
Due to the noisy and complex nature of global supply chains and firm-level transaction data, we completed additional processing of the data to support our analysis. This involved converting historical transactions to use the most recent version of the Harmonised System for product classification (HS 2022), establishing firm ownership structures to reconcile subsidiaries at the country level, and addressing the set of firms that do not complete a transformation on the goods they import, or their imports are unrelated to their exports.
HS 2022
As some transactions in our data use the previous 2017 Harmonised System of product classifications we convert these to the 2022 Harmonised System so that we have a consistent set. This was necessary as part of the period under study was in 2021 where the HS 2017 convention was still in place, and there were also cases where firms in later years had not updated their codes to the newer standard when completing customs declarations. We followed guidance produced by the Statistical Division of the UN for the conversion of products between the different harmonised systems which applies a matching process based on correlation tables.
Firm ownership
Within the data, transacting firms are reported at various levels within a company ownership structure; some may be wholly owned, while others might be a subsidiary of a multinational (or a subsidiary of a subsidiary … of a multinational).
The reconciliation of subsidiaries within country is important as the underlying data has instances whereby a subsidiary of a multi-national imports a product which is then used by a separate subsidiary of the same multi-national. In this case the imported goods are transferred between the subsidiaries domestically (under the same owner umbrella), but we do not capture these transfers as our data contains only cross-border trade.
Using GSCIP ownership data, we complete a process to agglomerate all firms to an ’owner-country’ level.
To rationalise the inconsistent ownership hierarchy, we use GSCIP’s ownership hierarchy dataset which sets out (where available) the global ultimate owner of subsidiaries. For each firm-entity we climb the ownership hierarchical structure until we get to the highest level of ownership up to the point of operation (as we do not want to classify distinct firms i.e. Heinz and Duracell into a larger non-distinct owner such as Berkshire and Hathaway.) To do this we stop climbing the ownership hierarchy to a higher-level firm when that firm can be classified as a financial corporation.
We also faced the issue of circular ownership - whereby a firm is owned by another one or more firms that are (at some tier) owned by the original firm - creating an infinite loop of ownership. To address this problem we break the circular ownership choosing one of the firms within the loop.
The ownership information we rely on is non-uniform and relationships can be opaque. This process improved the information we have on the firm set - but is however not a perfect process.
Once we have assigned subsidiaries to larger conglomerates we create a country-based version of each firm. So that, for example, a multinational firm that transacts to and/or from the UK would be treated as distinct to the same multinational that transacts to and/or from the US.
Additional analysis
Figure S1: [A-B] In and Out Degree. [C] Capital, intermediate and consumption goods (UN definition).
Community detection. We present the number of communities, the Stability coefficient and the Variation of Information for the SC and AI networks for each partition in Figure S2. As the partition number increases, the algorithm runs for a longer ’time’, and larger (but typically fewer) communities are detected.
At each scale (partition number), the algorithm seeks to minimise the Stability coefficient and hence the communities are better defined for lower values. In addition, a low value of the Variation of Information tells us we find similar partitions for different random seeds. We observe that both the SC and AI network have very well defined communities at large partition numbers (low Stability and low Variation of Information), and hence we focus on communities in this region. In Figure S2 C-E, we visualise partitions 20, 25 and 30.
In Figure S3 we show two partitions of the NAFTA IO network, which exhibit similarities to the SC and AI networks.
Figure S2: Community detection statistics: Number of communities, stability and variation of information for SC and AI networks.
Figure S3: IO communities: Two partitions of the Nafta IO network, which exhibit similarities to the SC and AI networks.
10 year growth regressions. In Figure S4 we replicate Figure 7 C and F for 10 year growth regressions (2011-2021). We find similar patterns with a slightly lower AUC in both cases.
Figure S4: 10 year growth regressions: